OpenStudy (anonymous):

A theater group made appearances in two cities. The hotel charge before tax in the second city was $1500 lower than in the first. The tax in the first city was 3%, and the tax in the second city was 9% . The total hotel tax paid for the two cities was $495.How much was the hotel charge in each city before tax? Please help!!

OpenStudy (anonymous):

@Talos could you help me with this one?

OpenStudy (anonymous):

Yeah, we need to take it step by step

OpenStudy (anonymous):

so x will be city 1

OpenStudy (anonymous):

y will be city 2

OpenStudy (anonymous):

first is the charge without tax so x=y+1500

OpenStudy (anonymous):

and the taxes equation will be .03x+.09y=495

OpenStudy (anonymous):

now substitute y=1500 for x in the second equation

OpenStudy (anonymous):

.0.3x+0.9y(1500)=495

OpenStudy (anonymous):

so now you have .03(y+1500) + .09y=495

OpenStudy (anonymous):

no you are substituting it for x

OpenStudy (anonymous):

because x=y+1500

OpenStudy (anonymous):

ooh, sorry

OpenStudy (anonymous):

Do you understand how I got there though?

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

ok now simplify for x

OpenStudy (anonymous):

How do I do that? o., there is no x in the equation

OpenStudy (anonymous):

Typo, I meant y

OpenStudy (anonymous):

Sorry about that

OpenStudy (anonymous):

are you gonna get 0.6y+1500=495?

OpenStudy (anonymous):

@Talos

OpenStudy (anonymous):

Not quite, do you know how the distributive property works?

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

Great! So distribute .03 over (y+1500)

OpenStudy (anonymous):

.03y+45

OpenStudy (anonymous):

yep now add .09y to that

OpenStudy (anonymous):

45.09

OpenStudy (anonymous):

or is it gonna be .03y+45+.09

OpenStudy (anonymous):

nope, notice the y so it would be .12y+45=495

OpenStudy (anonymous):

the y on .09y + .03y= .12y

OpenStudy (anonymous):

oh okay, i understand how u get that

OpenStudy (anonymous):

so now subtract 45 from both sides and get .12y=450

OpenStudy (anonymous):

divide both sides by .12 and you get?

OpenStudy (anonymous):

I got 3750

OpenStudy (anonymous):

yep y=3750, going back to where we started we know y is Hotel 2, so now we need the rate for Hotel 1

OpenStudy (anonymous):

use x=y+1500

OpenStudy (anonymous):

plug in 3750 for y

OpenStudy (anonymous):

x=3750+1500

OpenStudy (anonymous):

so x=5250

OpenStudy (anonymous):

SO your final answer is Hotel 1's rate: $5250 Hotel 2's rate: $3750

OpenStudy (anonymous):

Thanks very much :) I wrote tht all down incase I get another problem like that lol

OpenStudy (anonymous):

No problem! With these types of problems, make two equations and substitute(: